# Condensing and Extracting Against Online Adversaries

### v2

@@ 9.1 One Bit per Round
We will use the following protocol from [alonnoar93collective]:
@@
Our protocol will have two stages. In the first stage, we will use the lightest bin protocol from [feige99] until the number of players is small enough, and then in the second stage we use the protocol from Lemma 9.2. Let $C_{0}$ be a large constant that we set later. In particular, our final protocol will be:
@@
For $11\leq i\leq r+$ 1, let $g_{i}$ be the number of good players in $P_{i}$ and let $p_{i}=\left\lvert P_{i}\right\rvert$. As we always choose the lightest bin at each stage, $p_{i+1}\leq p_{i}/2$. Hence, we infer that $p_{i}\leq 2^{-i+1}\cdot\ell$. Let $g_{1}=g$. We next lower bound $g_{i}$:
@@
With probability at least $1-\exp(-(1/10)\cdot(g/2^{r}))$, it holds that for all $11\leq i\leq r+$ 1, $g_{i}\geq\frac{g}{2^{i}}-5\left(\frac{g}{2^{i}}\right)^{2/3}$.
@@ 9.2 Multiple Bits per Round
Our protocol and proof is similar to Lemma 9.1 with the key difference being that the larger value of $n$ allows us to increase the number of bins and simplify our analysis. Here, we end up being verbose and repeating ourselves for clarity. Just like earlier, our protocol will have two stages, one using the lightest bin protocol from [feige99] until the number of players is small enough and then resorting to the protocol from Lemma 9.2. Let $C_{0},C_{1}$ be large constants that we set later. Our final protocol will be:
@@
For $11\leq i\leq r+$ 1, let $g_{i}$ be the number of good players in $P_{i}$ and let $p_{i}=\left\lvert P_{i}\right\rvert$. As we always choose the lightest bin at each stage, $p_{i+1}\leq p_{i}/b_{i}$. Hence, we infer that $p_{r+1}\leq\ell/\prod_{i=1}^{r}b_{i}$. Let $g=g_{1}$. We first bound $g_{i}$:
@@
For any constant $C_{1}$, with probability at least $1-\exp(-\log(p_{r+1})^{1/5})$, it holds that for all $11\leq i\leq r+$ 1, $g_{i}\geq\frac{g}{\prod_{j=1}^{i-1}b_{j}}-2\left(\frac{g}{\prod_{j=1}^{i-1}b_{j}}\right)^{2/3}$.
@@ 10 Open Problems
  • Some of our condenser results are only existential and not explicit. It would be very interesting to find explicit constructions with similar parameters. As we show, one way of achieving this would be to explicitly construct a seeded condenser with dependence on seed length being $1\cdot\log(1/\varepsilon)$.
@@ Acknowledgements
We thank Madhur Tulsiani for asking a question that motivated us to consider the model of local oNOSF sources in Appendix A. We thank the organizers of the Dagstuhl Seminar on Algebraic and Analytic Methods in Computational Complexity and Schloss Dagstuhl for providing a stimulating research environment, where discussions between R.S. and E.C. contributed to this collaboration.
[…]
Appendix A Extracting from Local oNOSF Sources
@@
Definition A.1 (Local oNOSF sources ).
@@
Interestingly, weakening the adversary in this way converts our oNOSF source into a small-space source. These sources were first studied by [kamp_deterministic_2011] and we refer the reader to them for a definition and background. Since the adversarial blocks of an $s$-local ${\left(g,\ell,n,k\right)}$-oNOSF source can only depend on the binary string of length at most $sn$ to its left, we easily see that an $s$-local ${\left(g,\ell,n,k\right)}$-oNOSF source is samplable by a space-$sn$ source.
[…]
Theorem A.2 (Using the explicit extractor of [chattopadhyay_extractors_2022]).
@@
A similar result with slightly better entropy requirement, but constant error, can be obtained using the small-space extractor from [li2023two].

### v3

@@ 9.1 One Bit per Round
We will use the following protocol from [AN93]:
@@
Our protocol will have two stages. In the first stage, we will use the lightest bin protocol from [Fei99] until the number of players is small enough, and then in the second stage we use the protocol from Lemma 9.2. Let $C_{0}$ be a large constant that we set later. In particular, our final protocol will be:
@@
For $1\leq i\leq r+1$, let $g_{i}$ be the number of good players in $P_{i}$ and let $p_{i}=\left\lvert P_{i}\right\rvert$. As we always choose the lightest bin at each stage, $p_{i+1}\leq p_{i}/2$. Hence, we infer that $p_{i}\leq 2^{-i+1}\cdot\ell$. Let $g_{1}=g$. We next lower bound $g_{i}$:
@@
With probability at least $1-\exp(-(1/10)\cdot(g/2^{r}))$, it holds that for all $1\leq i\leq r+1$, $g_{i}\geq\frac{g}{2^{i}}-5\left(\frac{g}{2^{i}}\right)^{2/3}$.
@@ 9.2 Multiple Bits per Round
Our protocol and proof is similar to Lemma 9.1 with the key difference being that the larger value of $n$ allows us to increase the number of bins and simplify our analysis. Here, we end up being verbose and repeating ourselves for clarity. Just like earlier, our protocol will have two stages, one using the lightest bin protocol from [Fei99] until the number of players is small enough and then resorting to the protocol from Lemma 9.2. Let $C_{0},C_{1}$ be large constants that we set later. Our final protocol will be:
@@
For $1\leq i\leq r+1$, let $g_{i}$ be the number of good players in $P_{i}$ and let $p_{i}=\left\lvert P_{i}\right\rvert$. As we always choose the lightest bin at each stage, $p_{i+1}\leq p_{i}/b_{i}$. Hence, we infer that $p_{r+1}\leq\ell/\prod_{i=1}^{r}b_{i}$. Let $g=g_{1}$. We first bound $g_{i}$:
@@
For any constant $C_{1}$, with probability at least $1-\exp(-\log(p_{r+1})^{1/5})$, it holds that for all $1\leq i\leq r+1$, $g_{i}\geq\frac{g}{\prod_{j=1}^{i-1}b_{j}}-2\left(\frac{g}{\prod_{j=1}^{i-1}b_{j}}\right)^{2/3}$.
@@
10 Online Influence and Extraction Lower Bounds
Towards proving lower bounds on the possiblity of extraction from oNOSF sources , we introduce a new, natural notion of influence of Boolean functions, which we call online influence. For simplicity, we first start by considering the class of oNOBF sources , which corresponds to uniform $\left(g,\ell,n=1\right)$}-oNOSF sources.
We believe this is an interesting new measure and is worth studying in its own right, and we refer the reader to Example 10.6 for a couple of interesting examples. For monotone functions (and more generally, unate functions), it is not hard to see that online influence equals the usual notion of influence (see Lemma 10.4 for a proof). Thus, to find interesting properties of online influence (compared to standard influence, Definition 10.1), one must look at non-monotone (in fact, non-unate) Boolean functions.
The following natural question arises towards our goal of proving extractor lower bounds: for a function $f$, what is the maximum online influence out of all $n$ bits? For the usual notion of influence, this question was resolved by the well-known theorem of [KKL88], who showed there always exists a bit with influence at least $\operatorname*{Var}(f)\cdot\Omega\left(\frac{\log\ell}{\ell}\right)$.
Organization
We formally define the notion for Boolean functions and discuss some basic properties in Section 10.1. We establish tight bounds on the online influence for general functions, including a Poincaré style inequality, in Section 10.2. We provide an example exhibiting a separation between maximum (standard) influence and online influence in Section 10.3. Finally, in Section 10.4, we extend the definition of online influence to subsets of coordinates (and functions from $\Sigma^{n}\rightarrow\{0,1\}^{m}$, for arbitrary alphabet $\Sigma$). This allows us to prove the required lower bounds on extraction (and condensing) from oNOSF sources .
Notation
For convenience, we introduce some notation that we use for the rest of this section. For any bit $b\in\{0,1\}$, let $e(b)=(-1)^{b}$. For any Boolean function $f:\{0,1\}^{\ell}\rightarrow\{0,1\}$, let $e(f)$ denote the function $e(f)(x)=(-1)^{f(x)}$.
10.1 Basic Properties
Definition 10.1 (Influence).
For a function $f:\{0,1\}^{\ell}\to\{0,1\}$, the influence of the $i$-th bit is
and the total influence is
Definition 10.2 (Online influence).
and the total online influence is
Remark 10.3.
Lemma 10.4.
Proof.
Using the monotonicity of $f$, note that for any $x\in\{0,1\}^{i-1}$ and any $y\in\{0,1\}^{{\ell}-i}$, $f(x,1,y)\geq f(x,0,y)$. Thus, $\operatorname{\mathbf{oI}}_{i}[f]=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1},y\sim\mathbf{U}_{{\ell}-i}}[f(x,1,y)-f(x,0,y)]=\operatorname{\mathbf{I}}_{i}(f)$. ∎
10.2 A Poincaré Inequality for Online Influence
Theorem 10.5.
Example 10.6.
The PARITY function on $\ell$} bits $\bigoplus_{\ell}:\{0,1\}^{\ell}\to\{0,1\}$ for $i\in[{\ell}-1]$ has online influence $\operatorname{\mathbf{oI}}_{i}[\bigoplus_{\ell}]=0$, while $\operatorname{\mathbf{oI}}_{\ell}[\bigoplus_{\ell}]=1$. Thus, PARITY meets the lower bound of Theorem 10.5. We note that this is starkly different from regular influence where $\operatorname{\mathbf{I}}_{i}[\bigoplus_{\ell}]=1$ for all $i$.
To prove Theorem 10.5, we will use Boolean Fourier analysis. For any $f:\{0,1\}^{n}\to\{0,1\}$, $e(f)$ has a unique Fourier expansion given by: $e(f(x))=\sum_{S\subseteq[{\ell}]}\widehat{f}(S)\chi_{S}(x)$, where $\chi_{S}(x)=(-1)^{\sum_{i\in S}x_{i}}$ and $\widehat{f}(S)=\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{\ell}}[e(f)(y)\chi_{S}(y)]$. [^14] For simplicity of notation, we use $\widehat{f}(S)$ for $\widehat{e(f)}(S)$. Also recall that $\widehat{f}(\emptyset)=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{n}}[e(f)(x)]$, $\operatorname*{Var}(e(f))=\sum_{S\subseteq[{\ell}],S\neq\emptyset}\widehat{f}(S)^{2}$, and for any $S\neq T$, $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[\chi_{S}(x)\chi_{T}(x)]=0$. For more background, we refer the reader to the excellent book by O’Donnell [ODo14].
Lemma 10.7.
We first derive Theorem 10.5 using Lemma 10.7.
Proof of Theorem 10.5.
We start with the lower bound. We have,
where the inequality uses Lemma 10.7.
The upper bound is easy to derive as well.
$=\sqrt{{\ell}\operatorname*{Var}(e(f))}.$
This completes the proof. ∎
Claim 10.8.
Assuming the above claim, let us prove Lemma 10.7. We supply the proof of 10.8 below.
Proof of Lemma 10.7.
We first prove the inequality $\operatorname{\mathbf{oI}}_{i}(f)\geq\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}$. Since for any $x\in\{0,1\}^{i-1}$ we have $\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,1})(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,0})(y)]\right\rvert=2\left\lvert\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)\right\rvert$ by 10.8, and the fact that $\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,b})(y)]$ is in [1,1][-1,1] for all $x\in\{0,1\}^{i-1},b\in\{0,1\}$, it follows that $\left|\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)\right|$ is in $[0,1]$.
Thus,
$=\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}$ $\text{(derived above).}\qed$
Proof of 10.8.
$e(f)(z)$ $=\sum_{S\subseteq[{\ell}]}\widehat{f}(S)\chi_{S}(z)$
$=\sum_{S\subseteq[{\ell}]}\widehat{f}(S)\chi_{S}(x,b,y)$
We also have that
$e(f)(z)$ $=e(f)(x,b,y)$
$=e(f|_{x,b})(y)$
Therefore, Equation 5 and Equation 6 allow us to conclude that
Thus, we have
[…]
$=\sum_{T\subseteq[i]}\widehat{f}(T)\chi_{T}(x,b)$
We now plug this in to our definition of $\operatorname{\mathbf{oI}}_{i}[f]$ in Equation 4 to get a simplified expression. Recalling the fact that for any $x\in\{0,1\}^{n}$, f(x)=(1e(f)(x))/2f(x)=(1-e(f)(x))/2, we have
10.3 A Tight Example for Maximum Online Influence
Definition 10.9 (Maximum influence).
Corollary 10.10.
Proof.
Definition 10.11.
Lemma 10.12.
Proof.
Compared with the result of [KKL88] that $\operatorname{\operatorname{\mathbf{I}}_{\max}}[f]\geq\operatorname*{Var}(f)\cdot\Omega\left(\frac{\log{\ell}}{{\ell}}\right)$, this exhibits a separation between maximum (standard) influence and the online influence (of balanced functions).
Lemma 10.13.
Proof.
10.4 Online Influence of Sets and Extraction Lower Bounds
Definition 10.14 (Online influence, Definition 8.1 restated).
For any function $f:\Sigma^{\ell}\to\{0,1\}$, and any $B\subset[{\ell}]$, where $B=\{i_{1}<i_{2}<\ldots<i_{k}\}$, define $\operatorname{\mathbf{oI}}_{B}(f)$ as follows: an online adversary $\mathcal{A}$ samples a distribution $\mathbf{X}$ in online manner. It starts by sampling the variables $x_{1},x_{2},\ldots,x_{i_{1}-1}$ independently and uniformly from $\Sigma$, then picking the value of $x_{i_{1}}$ depending on $x_{<i_{1}}$. Next, the variables $x_{i_{1}+1},\ldots,x_{i_{2}-1}$ are sampled independently and uniformly from $\Sigma$, and $\mathcal{A}$ sets the value of $x_{i_{2}}$ based on all variables set so far, and so on. Define the advantage of $\mathcal{A}$ to be $\textrm{adv}_{f,B}(\mathcal{A})=|\operatorname*{\mathbb{E}}[f(\mathbf{X})]-\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell})]|$. Then $\operatorname{\mathbf{oI}}_{B}(f)$ is defined to be $\max_{\mathcal{A}}\{\textrm{adv}_{f,B}(\mathcal{A})\}$, where the maximum is taken over all online adversaries $\mathcal{A}$ that control the bits in $B$.
Definition 10.15.
and the total online influence is
Lemma 10.16 (online-resilient functions yield extractors).
Proof.
[…]
$\leq\varepsilon_{1}+\varepsilon_{2},$
as claimed. ∎
Remark 10.17.
Remark 10.18.
Theorem 10.19.
Proof.
We now repeat this process $t$ times to get $B_{t}=\{i_{1},\dots,i_{t}\}$ until our goal is achieved. For general $t$, let $f_{t}=f|_{B_{t}}$ where $B_{t}=B_{t-1}\cup\{i_{t}\}$ and $i_{t}=\operatorname*{\arg\!\max}_{i\in[n]\setminus B_{t-1}}\{\operatorname{\mathbf{oI}}_{i}[f_{t-1}]\}$. At the (t1)(t-1)-th step, since we have not stopped, it means that $\operatorname*{\mathbb{E}}_{x}[f_{t-1}(x)=1]<\beta$, but we of course have $\operatorname*{\mathbb{E}}_{x}[f_{t-1}(x)=1]\geq\alpha$ as well. Thus, by Corollary 10.10, collecting $i_{t}$ as a bad bit gives us that
$=\alpha+\frac{4\alpha(1-\beta)}{{\ell}}\cdot t.$
Corollary 10.20.
Proof.
Corollary 10.21.
For any balanced function $f:\{0,1\}^{\ell}\to\{0,1\}$ and $0<\varepsilon<1/3$, there exists a $\left(g=\ell-b,\ell\right)$}-oNOBF source $\mathbf{X}$ with $b\leq 3\varepsilon\ell$ such that $\left\lvert f(\mathbf{X})-\mathbf{U}_{1}\right\rvert\geq\varepsilon$.
Proof.
Remark 10.22.
@@ 11 Open Problems
  • While we obtain explicit condensers for almost all parameter regimes, it remains open to construct them when the bock length is constant, matching the parameters of our existential results. As we show, one way of achieving this would be to explicitly construct a seeded condenser with dependence on seed length being $1\cdot\log(1/\varepsilon)$.
@@ Acknowledgements
We thank Madhur Tulsiani for asking a question that motivated us to consider the model of local oNOSF sources in Appendix B. We thank the organizers of the Dagstuhl Seminar on Algebraic and Analytic Methods in Computational Complexity and Schloss Dagstuhl for providing a stimulating research environment, where discussions between R.S. and E.C. contributed to this collaboration.
Appendix A Constructing $\operatorname{\mathsf{Reduce}}^{\prime}$
In this section we construct $\operatorname{\mathsf{Reduce}}^{\prime}$ which has the properties as guaranteed by Lemma 5.12. In [RZ01], the authors use hitting sets for combinatorial rectangles to reduce $\ell$-length oNOSF sources to shorter min-entropy sources. We provide a proof of their lemma for completeness here.
Let’s first define combinatorial rectangles.
Definition A.1 (Combinatorial rectangle).
Let $a,d\in\mathbb{N}$. We say that a set $R\subseteq[a]^{d}$ is a combinatorial rectangle if $R=R_{1}\times R_{2}\times\cdots\times R_{d}$ for some sets $R_{i}\subseteq[a]$ for $i\in[d]$. The density of $R$ is $\operatorname{Density}(R)=\frac{1}{a^{d}}\prod_{i=1}^{d}\left\lvert R_{d}\right\rvert$.
A hitting set for a family of combinatorial rectangles is a subset of $a]^{d}$ such that it has an intersection with every combinatorial rectangle in the family. Formally:
Definition A.2 (Hitting sets for combinatorial rectangles).
A set $\mathcal{H}\subseteq[a]^{d}$ is a $(a,d,\delta)$-hitting set for combinatorial rectangles if for every combinatorial rectangle $R\subseteq[a]^{d}$ with $\operatorname{Density}(R)\geq\delta$ we have that $R\cap\mathcal{H}\neq\varnothing$.
Of course, taking $\mathcal{H}=[a]^{d}$ is a trivial hitting set for any combinatorial rectangle, so the difficulty lies in decreasing the cardinality of $\mathcal{H}$ while keeping the density requirement $\delta$ of the combinatorial rectangle low. In [LLSZ97], the authors create a small enough hitting set for our use.
Lemma A.3 ([LLSZ97]).
There exists a universal constant $C$ such that for any $\delta>0$ and $a,d\in\mathbb{N}$, there exists an explicit construction of an $(a,d,\delta)$-hitting set $\mathcal{H}\subseteq[a]^{d}$ such that $\left\lvert\mathcal{H}\right\rvert\leq\left(\frac{a\log(d)}{\delta}\right)^{C}$.
Let’s see how using all of these ingredients we can construct $\operatorname{\mathsf{Reduce}}^{\prime}$.
Proof of Lemma 5.12.
To construct $\operatorname{\mathsf{Reduce}}^{\prime}$, we begin by defining a family of functions $\mathcal{F}\subseteq\{f:[a]^{d}\to\{0,1\}\}$ and a hitting set $\mathcal{H}\subseteq[a]^{d}$ of size $\left\lvert\mathcal{H}\right\rvert=2^{t}=T$. For every $x\in\operatorname{Supp}(\mathbf{X})$, we will select a $f_{x}\in\mathcal{F}$ and output the smallest $y\in\mathcal{H}$ such that $f_{x}(y)=1$, where we consider our output as an element of $\left\lvert\mathcal{H}\right\rvert]=\{0,1\}^{t}$. Then, for all $y\in\mathcal{H}$, we will show that $\Pr_{x\sim\mathbf{X}}[f_{x}(y)=1]\leq 2^{-k}$, meaning that $\operatorname{\mathsf{Reduce}}^{\prime}(\mathbf{X})$ is a $(t,k)$-source.
Formally, we let $\mathcal{F}$ be the following family of combinatorial rectangles on $a]^{d}$. Given an $x\in[a]^{d}$, we define the combinatorial rectangle $\operatorname{\mathsf{Rect}}_{x}=\{y\in[a]^{d}\mid\forall i\in[d],y_{i}\neq x_{i}\}$ and the associated function $f_{x}:[a]^{d}\to\{0,1\}$ for this rectangle as $f_{x}(y)=1_{y\in\operatorname{\mathsf{Rect}}_{x}}$. Then, we let $\mathcal{F}=\{f_{x}\mid x\in[a]^{d}\}$.
Note that the density of any particular rectangle $\operatorname{\mathsf{Rect}}_{x}$ is $\delta=\operatorname{Density}(\operatorname{\mathsf{Rect}}_{x})=\frac{\operatorname{\mathsf{Rect}}_{x}}{a^{d}}=\left(1-\frac{1}{a}\right)^{d}$. We can lower bound $\delta$ as $\delta\geq(\exp(-d/a))^{C_{\delta}}$ for some universal constant $C_{\delta}$. Rearranging then gives us that $\log(1/\delta)\leq C_{\delta}\cdot\frac{d}{a}$. With this in mind, we set up our hitting set $\mathcal{H}$ for $\mathcal{F}$. From [LLSZ97], we know that there exists a universal constant $C_{1}$ and an explicit hitting set $\mathcal{H}$ such that $T=\left\lvert\mathcal{H}\right\rvert\leq\left(\frac{a\log(d)}{\delta}\right)^{C_{1}}$. Simplifying this expression yields
$T$ $\leq\left(\frac{a\log(d)}{\delta}\right)^{C_{1}}$
$t$ $\leq C_{1}(\log(a)+\log\log(d)+\log(1/\delta))$
$t$ $\leq C_{1}(\log(a)+\log\log(d)+C_{\delta}d/a)$
$t$ $\leq C^{\prime}(\log(a)+\log\log(d)+d/a),$ (7)
where $C^{\prime}$ is a sufficiently large universal constant, depending only on $C_{1},C_{\delta}$. To analyze the min-entropy of $\operatorname{\mathsf{Reduce}}^{\prime}(\mathbf{X})$, we note that for all $y\in[a]^{d}$
$\Pr_{x\sim\mathbf{X}}[f_{x}(y)=1]$ $\leq\left(1-\frac{1}{a}\right)^{\gamma d}$
$\leq\exp(-\gamma d/a)\leq 2^{-\gamma d/a},$
which directly implies that the min-entropy $k$ of $\operatorname{\mathsf{Reduce}}^{\prime}(\mathbf{X})$ is $\geq\gamma d/a$, as desired. ∎
Appendix B Extracting from Local oNOSF Sources
@@
Definition B.1 (Local oNOSF sources ).
@@
Interestingly, weakening the adversary in this way converts our oNOSF source into a small-space source. These sources were first studied by [KRVZ11] and we refer the reader to them for a definition and background. Since the adversarial blocks of an $s$-local $\left(g,\ell,n,k\right)$}-oNOSF source can only depend on the binary string of length at most $sn$ to its left, we easily see that an $s$-local $\left(g,\ell,n,k\right)$}-oNOSF source is samplable by a space-$sn$ source.
Using recent explicit extractors for low-space sources provided by [CL22, Li23] and the fact that a ${\left(g,\ell,n,k\right)}\textrm{-oNOSF source}$ has entropy at least $gk$, we get the following extraction result for these local online sources.
Theorem B.2 (Using the explicit extractor of [CL22]).
@@
A similar result with slightly better entropy requirement, but constant error, can be obtained using the small-space extractor from [Li23].